23 (6x−9)=1/2 (7x+6)
Answer by stanbon(75887)

(6x−9)=1/2 (7x+6)
----
2(6x-9) = 7x+6
--------------------
12x - 18 = 7x + 6
5x = 24
x = 4.8
-------------
Cheers,
Stan H.
-----------